OpenStudy (anonymous):

solve the system of equations {x^2 - 3xy + 2y^2 = 0 and x^2 + y^2 = 0} can yuu show me how to work out this problem?

OpenStudy (anonymous):

try taking the second equation and solve for y, then substitute this value into the first equation

OpenStudy (anonymous):

Note that you can rewrite the first equation as:\[x^2-3xy+2y^2=0\Longrightarrow (x^2+y^2)-3xy+y^2=0\]However, you know that:\[x^2+y^2=0\]So now we obtain:\[(x^2+y^2)-3xy+y^2=0\Longrightarrow -3xy+y^2=0\]This will be easier to work with for solutions.